. Blue Enterprises paid a Birr 2.50 per share dividend on 100,000 shares of common stock and Birr 4 per share dividend on 25,000 shares of preferred stock. Net income for the year was Birr 850,000. The current market price of the common stock is Birr 52.50 a) Calculate the earnings per share of common stock. c) Calculate the dividend payout ratio of common stock d) Calculate the dividend yield of common stock


1
Expert's answer
2022-01-02T18:20:14-0500

Solution:

a.). EPS of common stock = (Net income – Preferred Dividends) "\div" Common share outstanding

Preferred Dividends = 25,000 shares "\times" 4 = Birr 100,000

= (850,000 – 100,000) "\div" 100,000 = 750,000 "\div" 100,000 = 7.5

EPS of common stock = Birr 7.5

 

c.). Dividend payout ratio of common stock = Total dividends "\div" Net income x 100

Total dividends to common stock = 100,000 "\times" 2.5 per share = 250,000

Dividend payout ratio of common stock = (250,000 "\div" 850,000) "\times" 100 = 29.41%

 

d.). Dividend yield of common stock = Annual dividends paid per share "\div" Pricr per share

Annual dividends paid per share = 2.50

Price per share/market value per share = 52.50

Dividend yield of common stock = 2.50 "\div" 52.50 = 0.048